### v1

# SpaCE-10: A Comprehensive Benchmark for Multimodal Large Language Models in Compositional Spatial Intelligence

Abstract
Multimodal Large Language Models (MLLMs) have achieved remarkable progress in various multimodal tasks. To pursue higher intelligence in space, MLLMs require integrating multiple atomic spatial capabilities to handle complex and dynamic tasks. However, existing benchmarks struggle to comprehensively evaluate the spatial intelligence of common MLLMs from the atomic level to the compositional level. To fill this gap, we present SpaCE-10, a comprehensive benchmark for compositional spatial evaluations. In SpaCE-10, we define 10 atomic spatial capabilities, which are combined to form 8 compositional capabilities. Based on these definitions, we propose a novel hierarchical annotation pipeline to generate high-quality and diverse question-answer (QA) pairs. With over 150+ hours of human expert effort, we obtain over 5k QA pairs for 811 real indoor scenes in SpaCE-10, which covers various evaluation settings like point cloud input and multi-choice QA. We conduct an extensive evaluation of common MLLMs on SpaCE-10 and find that even the most advanced MLLM still lags behind humans by large margins. Through our careful study, we also draw several significant findings that benefit the MLLM community. For example, we reveal that the shortcoming of counting capability greatly limits the compositional spatial capabilities of existing MLLMs. The evaluation code and benchmark datasets are available at here.
Figure 1: Overview of SpaCE-10 benchmark. SpaCE-10 takes over 150 human expert hours to collect 5k+ QA pairs in 811 indoor scenes, which can evaluate MLLMs from 10 atomic capabilities to 8 compositional capabilities. Through extensive evaluations, SpaCE-10 indicates that even the most advanced MLLM still lags far behind humans by large margins.
1 Introduction
Recent years have witnessed the rapid development of multimodal large language models (MLLMs) gpt4 ; gemini ; claude3 ; internvl ; qwen ; blip ; blip2 , which continually narrows the gap between machines and humans in various multimodal tasks mmbench ; okvqa ; gqa ; coco_caption ; evocodebench . The significant progress has motivated researchers to pursue higher machine intelligence in the real world leo ; 3d-llm ; gpt4scene ; chat-scene ; pointllm ; guo2023point ; rovi ; vebrain ; interleave-vla . As shown in Fig. 1, when situated in a real scene, the machine needs to perceive the surrounding environment, collect key content, analyze visual cues, and perform interactions. This places a higher demand on existing MLLMs in compositional spatial intelligence, where multiple atomic spatial capabilities should be integrated together in task solving. Therefore, a question is naturally raised: Can existing MLLMs master atomic space capabilities and seamlessly combine them?
Nevertheless, this question still remains unanswerable via existing evaluation benchmarks 3dqa ; 3dsrbench ; scanqa ; sqa3d ; clever3d ; msqa . As shown in Tab. 1, early benchmarks scanqa ; sqa3d ; clever3d ; 3dqa ; scanrefer mainly focus on the assessment of atomic capabilities like object recognition and spatial localization, while compositional ones still remain to be defined and evaluated. Recently, VSI-Bench vsi aims to evaluate the visual-spatial intelligence of MLLMs through more compositional questions, but still fails to reflect the role of different atomic capabilities in compositional reasoning. More importantly, existing spatial benchmarks struggle to satisfy the evaluation needs of existing MLLMs in terms of scenes, modalities, and question types, etc. As shown in Tab. 1, the number of scenarios in existing benchmarks is usually less than 400, which makes it difficult to cover various practical situations.
To fill these gaps, we present SpaCE-10, a comprehensive benchmark for compositional Spatial Capability Evaluation of MLLMs. In SpaCE-10, we first define 10 atomic spatial capabilities that common MLLMs should master in the real world, e.g., object recognition and spatial localization. Based on these atomic capabilities, we construct a pool of compositional capabilities, each of which requires combining multiple atomic ones. As shown in Fig. 1, one question-answer pair often consists of more than 3 atomic capabilities. Therefore, SpaCE-10 can not only assess the compositional spatial capabilities of MLLMs, but can also reflect the impact of different atomic capabilities in compositional reasoning.
Based on this design principle, we propose an innovative hierarchical annotation pipeline in SpaCE-10. Specifically, we collect over 800 real indoor scanned scenes from four public datasets. For each scene, we present an automated pipeline to generate structured data that can describe different types of information in the scene, e.g., appearance and relationship. Based on this information, a multi-stage semi-automated pipeline is adopted to generate basic QA pairs, conduct quality verification, and perform the capability integration. Our SpaCE-10 consists of more than 5,000 high-quality QA pairs, covering various settings of existing MLLMs, e.g., point cloud input and multi-choice question types. As shown in Tab. 1, SpaCE-10 demonstrates greater diversity than previous benchmarks in data distribution, annotation process, and evaluation settings, showing promising all-around evaluation ability for compositional spatial intelligence.
We conduct extensive and systematic evaluations of mainstream MLLMs on SpaCE-10, including 3 close-source MLLMs and more than 20 open-source MLLMs ranging from 2B to 78B. Experimental results show that even the most advanced MLLMs are still far behind humans in compositional spatial intelligence, i.e., 37.4% of GPT4o vs. 72.0% of human. Meanwhile, 2D MLLMs demonstrate much stronger capabilities than 3D MLLMs on SpaCE-10, showing great potential for image-based spatial reasoning. Besides, existing MLLMs greatly fall short in multiple-choice QAs, suggesting their inferior complex reasoning abilities. Our further study also reveals that the shortcoming of counting capability greatly limits the compositional spatial capabilities of existing MLLMs. These findings provide valuable directions for the community to develop more capable MLLMs. Overall, our main contributions are summarized as follows:
  • We present SpaCE-10, a comprehensive benchmark for compositional spatial intelligence. SpaCE-10 is the most diverse benchmark that can assess the capabilities of MLLMs from the atomic level to the compositional level. It also covers various evaluation settings, including 3D inputs and multi-choice questions.
  • We propose an innovative hierarchical annotation pipeline in SpaCE-10, which first produces structured descriptions of scenes via an automated pipeline and then generates compositional QA pairs through a multi-stage semi-automated pipeline. The hierarchical pipeline ensures the quality, diversity, and controllability of the generated QA pairs.
  • We conduct extensive evaluations for nearly 30 open- and close-source MLLMs on SpaCE-10. Through our in-depth analysis, we draw several significant findings that will benefit the deployment of future MLLMs in the community.
Table 1: Comparison of SpaCE-10 with existing spatial benchmarks. Our SpaCE-10 contains the most diverse scenarios and QA types, covering various evaluation settings of existing MLLMs.
Dataset Scenario Source Scene Q&A Metric 2D & 3D Multi-Choice Multi-Capability
3DQA fe-3dgqa SCN scannet - 902 Sim. $\times$× $\times$× $\times$×
ScanQA scanqa SCN scannet 167 10k Sim. $\checkmark$ $\times$× $\times$×
FE-3DGQA fe-3dgqa SCN scannet 100 3.9k Sim. $\times$× $\times$× $\times$×
SQA3D sqa3d SCN scannet 132 6.9k Sim. $\checkmark$ $\times$× $\times$×
CLEVER3D clever3d SCN scannet 133 10k Sim. $\times$× $\times$× $\times$×
3D-LLM 3d-llm OBJ,SCN scannet ,HM3D hm3d - 30k Sim. $\times$× $\times$× $\times$×
M3DBench m3dbench SCN scannet - 1.5k Sim.+LLM $\times$× $\times$× $\times$×
MSQA msqa SCN scannet ,3RS 3rscan ,ARK arkitscenes 381 3.5k LLM $\times$× $\times$× $\times$×
VSI vsi SCN scannet ,3RS 3rscan ,SCN++ scannet++ 288 5.0k Acc. $\times$× $\times$× $\checkmark$
SpaCE-10 (Ours) SCN scannet ,3RS 3rscan ,ARK arkitscenes ,SCN++ scannet++ 811 6k Acc. $\checkmark$ $\checkmark$ $\checkmark$
2 Related Work
2.1 Multimodal Large Language Models
Advancements in large language models (LLMs) have led to the development of multimodal large language models (MLLMs) like GPT gpt4 , Claude claude3 , Gemini gemini ; gemini1.5 , and open-source models such as BLIP blip ; blip2 ; instructblip , LLaVA llava-v1.5 ; llavanext ; llava-onevision , Qwen qwen ; qwen2.5 ; qwen2.5-vl ; qwen2.5-math , and InternVL internvl ; internvl2 ; internvl2.5 ; internvl3 ; monointernvl . However, studies eye_wide_shut ; 3dsrbench ; vsi reveal that MLLMs still struggle with visual spatial intelligence, particularly in understanding object relationships, positions, and maintaining consistency across viewpoints. Recent efforts pointllm ; scene-llm ; leo ; llava-3d ; gpt4scene to integrate 3D point clouds data with 2D images have improved spatial reasoning but still face challenges in aligning geometric and semantic information, limiting their real-world applicability.
2.2 Multimodal QA Benchmarks
With the expanding capabilities of MLLMs, an increasing number of studies have focused on systematically evaluating their performance across diverse tasks, including caption generation coco_caption ; nocaps ; young2014image , visual question answering (VQA) gqa ; okvqa ; vqav2 ; visual7w ; mminstruct ; mmbench ; scienceqa ; ocrbench , and code generation evocodebench . Among these tasks, VQA has emerged as one of the most widely adopted formats due to its alignment with human interaction paradigms. Despite strong performance on general VQA tasks, assessing MLLMs’ spatial intelligence remains a major challenge. Existing benchmarks in this area fall into two main categories. The first logicvista ; visulogic focuses on 2D abstract reasoning, using logic puzzles or geometric patterns to test panel spatial reasoning. The second 3dsrbench ; scanqa ; sqa3d ; visual_spatial_reasoning ; vsi ; omnispatial ; mmsi ; ma_survey ; spartun3d ; space3d explores realistic scene understanding, using natural images or 3D point clouds data. Our SpaCE-10 begins from the capability and evaluates MLLMs’ compositional spatial intelligence.
3 SpaCE-10
3.1 Data Collection
Figure 2: Statistics of SpaCE-10. (Left) Distribution of 8 compositional question types in SpaCE-10. (Middle) Proportion of questions across different scenarios. (Right) The 10 atomic capabilities evaluated in SpaCE-10 and their coverage across all questions.
SpaCE-10 comprises several 3D indoor-scene scan datasets, including ScanNet++(scannet++, ), ScanNet(scannet, ), 3RScan (3rscan, ), and ARkitScene (arkitscenes, ). For scene selection, we use the ScanQA scanqa test set for sampling scenes from ScanNet, and select scenes from the validation and test sets of the other scan datasets. To support both 2D and 3D data inputs, we collect the official 2D frames corresponding to the 3D scans from each dataset. The final dataset consists of over 800 real indoor scenes, including a wide variety of indoor environments such as living rooms, classrooms, bathrooms, kitchens, and more. The dataset also includes more than 5,000 high-quality question-answer pairs. The number of QA pairs in each scan dataset is detailed in Figure 2 (Middle).
3.2 Hierarchical Annotation Pipeline
Overview. As shown in Fig 3 (a), our annotation pipeline consists of 5 stages from data preparation to high-quality QA generation. In stage-1, we employ human experts to manually collect snapshots of 3D point cloud scans from 4 to 6 different views, which costs over 38 hours of human expertise to check the quality of snapshots. In stage-2, we combine the collected snapshots and video frames to generate the structural data of describes different information in the scene. Then, we use this data to generate over 10k basic QAs covering atomic capabilities. In stage-4, human experts will manually filter the low-quality QA pairs, costing over 112 expert hours and resulting in over 8k+ QA pairs. Finally, we design a cross-capability integration strategy to mix QAs of different atomic capabilities into a single one, yielding the final QA pairs covering 8 compositional spatial capabilities.
Figure 3: Illustration of our hierarchical annotation pipeline. We generate structural data to construct over 10k QA pairs, and performs capability integration to obtain over 5k QA pairs with 10 compositional capabilities. This process takes over 150 expert hours for data collection and filtering.
Structural Data Generation.
As shown in Fig. 3 (b), this pipeline follows a progressive design with 6 steps: (1) In step-1, we select 10 keyframes from the video of each scenario by combining the CLIP vision encoder clip and the k-means algorithm. (2) In step-2, we leverage GPT-4o to generate a 2D caption for each scene, which covers information of appearance, size, and spatial relationships. (3) In step-3, we reuse GPT-4o as an inspector to refine the 2D captions by removing incorrect and redundant information. (4) In step-4, our manually collected 3D snapshots will be combined with keyframes for 3D caption generation. These high-quality snapshots contain rich global information about the whole scenes, which can provide considerable scene-level spatial information. (5) In step-5, we leverage an inspector again to further check and refine the spatial descriptions. (6) Finally, we employ a rule-based extractor to obtain structural data for the following QA generation.
[…]
During the basic QA generation, we adopt 3 approaches: template-based, MLLM-based, and human-based generation. For EQ, we first extract the number of each object from the scan datasets’ semantic labels, and then leverage GPT-4o to generate QA. For the SP, we employed human experts to manually design 80 QA pairs focused on tidying a room. Each question presents a complete task flow, including the navigation path, goal, goal characteristics, and actions to be performed, with potential errors in any step. Based on the prompt, the model must select either a fully correct task flow or one containing incorrect steps. For the rest of the QA types, we generate based on structural data by GPT-4o. Moreover, 5 types of QA will be applied with a cross-capability mixture, and we will introduce them next.
Cross-Capability Integration Strategy. After generating the basic QA, we proceed to enhance the capabilities of the five types of questions in the basic QA. Before introducing the enhancement, we briefly describe the atomic capabilities. SpaCE-10 establishes an atomic capability pool, including c1 to c10: Object Recognition, Spatial Localization, Spatial Relationship, Size Comparison, Counting, Function Knowledge, Multi-view Fusion, Forward Thinking, Reverse Reasoning, and Situated Observation. As shown in Fig. 2 (Right), in the capability pool, the most frequently utilized capabilities are C1, C2, C7, and C8, forming the spatial perception foundation of MLLMs. Higher-level abilities, such as C3, C9, and C10, are tested in approximately 50% of the questions, enhancing the model’s spatial reasoning ability. More details of capabilities are provided in the supplementary material.
Based on these atomic capabilities, we apply three strategies to integrate cross-capabilities into five types of QA: (1) For SA, OO, and OS, we integrate an additional C7. In the original basic QA, each option involves only a single object from the same scene. For OO and OS, we enhance them from single-object examinations to multiple objects from the same scene in each option, converting a single view to multiple-view fusion. Additionally, for OS, we slightly adjust the strategy by combining multiple spatial relationships into a single question, e.g., left, right, farthest, nearest, and so on. This integration forces the model to fully comprehend the entire scene, enabling a more holistic spatial perception. (2) For EP, we add C7 and C9, expanding the question to involve multiple objects. We then reverse the question type from “which object exists” to “which object does not exist,” integrating reverse reasoning capabilities. (3) For FR, we add C7, C9, and C10. The basic question asks which option correctly describes the function of an object near a given central object. The options are structured as <Object1>, <Function>…. In the integrated version, the question is revised to: “Which of the following is correct?” Each option now involves a different central object, with the structure: “Among the objects adjacent to <Central Object1>, one can be used for <Function>, while none of the others can.” This change prevents potential leakage of prior knowledge, such as object names, and places greater emphasis on the model’s understanding of functional roles. Similar to the previous strategy, the task scope is expanded to cover the entire scene.
3.3 Quality Verification
For the quality verification process, we rely on manual screening by human experts. We set up a validation framework based on Gradio and employed two human experts to perform the evaluation. In this process, the evaluation criteria include checking for incorrect options, invalid answers, missing data, or questions involving objects not present in the snapshots. This quality control process takes over 112 hours to check and filter low-quality QA pairs. By employing human validation, we ensure that only high-quality and contextually accurate questions are retained for the final benchmark. Related visualizations are attached in the supplementary material.
4 Experiments
4.1 Setup
In our experiments, we test extensive close-source and open-source MLLMs on SpaCE-10, including 3D MLLMs leo leo , MLLMs for 3D GPT4Scene gpt4scene , GPT-4o gpt4 , Gemini-2.0 gemini , Claude-3.7-sonnet claude3 , InternVL2.5 internvl2.5 , Qwen2.5-VL qwen2.5-vl , and so on. During the evaluation, except for Leo, which is evaluated by their own framework, other MLLMs are evaluated by the LMMs-Eval lmms-eval , and open-source MLLMs are tested on Tesla A100 GPUs. For the response and answer alignment, we follow the prompt of mmbench mmbench and use GPT-4o-2024-11-20 for the judgment. Notably, we removed the random assignment in LMMs-Eval, so the model’s performance is likely to be lower than the random baseline (25% for single choice, 10% for double choice).
4.2 Overall Results
Human vs. MLLMs. We first test human performance on SpaCE-10, taking the average score of two humans to obtain the ‘Human’ score in Tab. 2. The results indicate that although human performance does not meet expectations when facing with more complex numerical and reasoning tasks, the total score of 72.0% is still significantly higher than all existing MLLMs. In comparison, the best open-source MLLM only achieves an average score of 46.7%, and the best close-source models only achieve a score of 37.4%. These results demonstrate that the compositional spatial intelligence of MLLMs is far below the human level.
Table 2: Main results of existing MLLMs on SpaCE-10 benchmark. The * means these QA pairs have the multiple-choice version, and the results are the weighted sum of single-choice and multiple-choice Accuracy.
Models Rank Perception Reasoning Overall
EQ SQ SA $O$ OS EP FR SP
Human 1 63.6 78.9 70.6 68.3 82.2 81.8 68.2 62.5 72.0
3D MLLMs
LEO-7B leo 28 15.8 0.0 16.7 16.5 25.2 5.5 5.7 13.3 12.3
GPT4Scene-Qwen2-VL-7B gpt4scene 20 30.9 26.9 31.8 31.5 23.6 19.6 27.7 32.5 28.0
Close Source 2D MLLMs
GPT-4o-2024-11-20 gpt4 7 35.3 32.5 53.5 51.3 29.6 38.8 34.5 23.7 37.4
Gemini-2.0-Flash-Thinking-Exp gemini 14 30.4 26.4 46.8 34.3 21.0 25.6 33.8 31.2 31.1
Claude-3.7-Sonnet-20250219 claude3 9 36.3 35.9 41.1 37.4 22.7 25.6 42.6 25.0 33.3
Open Source 2D MLLMs
[…]
InternVL2.5-1B internvl2.5 22 33.0 40.1 18.3 30.3 17.4 26.3 23.7 23.7 26.6
InternVL2.5-2B internvl2.5 21 32.2 28.0 24.4 32.0 20.9 31.4 13.7 36.2 27.3
Qwen2.5-VL-3B-Instruct qwen2.5 16 31.7 45.5 39.0 43.0 25.3 11.5 22.8 21.2 30.0
VILA1.5-3B vila 24 34.0 9.1 23.7 26.5 22.4 10.3 32.1 33.7 23.9
InternVL2.5-4B internvl2.5 8 34.3 34.4 43.6 44.4 16.5 31.1 50.1 33.7 36.0
[…]
Qwen2.5-VL-7B-Instruct qwen2.5 17 32.7 31.0 41.3 32.1 27.6 15.4 23.6 27.5 28.9
LLaVA-v1.5-7B liu2023visual 13 31.2 31.3 30.5 35.7 22.9 10.7 57.4 32.5 31.5
LLaVA-OneVision-7B llava-onevision 10 37.4 36.2 42.9 44.4 27.1 11.2 45.6 21.2 33.2
InternLM-XComposer2.5-7B internlm-xcomposer 18 29.8 26.3 34.9 37.5 29.1 5.6 27.8 35.0 28.2
Cambrian-8B cambrian 23 22.6 18.6 27.2 27.6 24.1 21.4 37.2 23.7 25.3
VILA1.5-8B vila 25 25.7 8.21 27.5 32.7 17.2 12.4 26.7 23.7 21.7
InternVL2.5-8B internvl2.5 12 33.2 30.6 41.6 49.9 27.4 14.0 29.8 32.5 32.3
[…]
Cambrian-13B cambrian 26 20.6 23.8 20.8 27.1 23.0 12.9 27.8 15.0 21.3
CogVLM2-LLaMA3-Chat-19B cogvlm 6 32.4 43.7 37.5 42.6 26.3 40.1 56.4 30.0 38.6
[…]
InternVL2.5-26B internvl2.5 5 34.3 38.1 47.9 48.6 28.2 51.3 36.1 33.7 39.7
Cambrian-34B cambrian 27 17.5 21.0 24.6 21.9 19.0 18.7 19.0 17.5 19.9
Qwen2.5-VL-32B-Instruct qwen2.5 19 19.9 26.9 44.1 32.2 24.5 11.7 33.3 32.5 28.1
InternVL2.5-38B internvl2.5 4 38.1 39.5 51.0 50.6 27.1 50.5 50.1 37.5 43.0
VILA1.5-40B vila 15 34.7 16.7 47.7 53.1 34.2 10.4 29.9 21.2 30.9
[…]
LLaVA-OneVision-72B llava-onevision 2 44.1 29.9 56.4 55.0 32.8 60.1 59.7 36.2 46.7
Qwen2.5-VL-72B-Instruct qwen2.5 11 32.4 34.7 46.9 32.7 26.3 23.7 30.8 33.7 32.6
InternVL2.5-78B internvl2.5 3 27.8 40.0 52.8 51.6 39.4 47.1 54.7 40.0 44.1
3D MLLMs vs. 2D MLLMs. In Tab. 2, we evaluate LEO and GPT4Scene as representatives of 3D-related MLLMs. Notably, input to the LEO model requires point clouds of the objects relevant to the questions. To ensure a fair comparison, we adjust the input for LEO by randomly sampling 1024 points from the entire scene’s point clouds. The results show that LEO scored 12.3% overall in SpaCE-10, which is significantly lower than GPT4Scene (28.0%). Compared with InternVL2.5-1B, the performance of LEO, which has 7B parameters, is also substantially lower. We argue that one of the limitations of current 3D MLLMs is that they are designed to focus on specific objects and have difficulty processing the entire scene’s point clouds as input. Additionally, they likely sacrifice multimodal conversational abilities for understanding scans. These results also indicate that 2D MLLMs have greater potential in visual spatial intelligence comprehension than 3D MLLMs.
Open-Source vs. Close-Source. In close-souce MLLMs experiments, GPT-4o-2024-11-20 gpt4 achieves the best overall performance, ranking 7th with a score of 37.4%. It excels in perception tasks, particularly in SA with 53.5% and OO with 51.3%, indicating strong accuracy in recognizing size and position. Additionally, GPT-4o gpt4 achieves the highest score in EP among close-source models, with a score of 38.8%. However, it struggles with SP, where it scores the lowest among all tasks, suggesting a limitation in scene-level planning ability. Gemini-2.0-Flash-Thinking-exp gemini and Claude-3.7-Sonnet-20250219 claude3 achieve similar overall scores of 31.1% and 33.3%, respectively. Compared to these models, Claude-3.7-Sonnet shows a more balanced performance across tasks. In open-source MLLMs, among models with a scale of over 72B, LLaVA-OneVision-72B delivers outstanding performance, ranking first overall with a score of 46.7%. It outperforms other models in nearly all tasks, solidifying its top position. Furthermore, InternVL2.5-78B achieves the second-highest score of 44.1%, showing the increasing competitive edge of open-source models in spatial reasoning. These results suggest that the gap between open-source and close-source models has significantly narrowed, and open-source MLLMs even outperform close-source models, especially in compositional spatial intelligence.
Single-Choice vs. Multiple-Choice. The results in Tab. 3 show that MLLMs perform significantly worse on multiple-choice tasks compared to single-choice tasks. Smaller models like InternVL2.5-1B and 2B score over 30% on single-choice tasks, while in multiple-choice tasks, their scores often fall below 10%, which is even worse than random selection. This trend is also observed in larger models like Cambrian-8B, 13B, and 34B, which perform poorly in multiple-choice evaluations despite excelling in single-choice scores. As the model size increases to 72B, MLLMs show greater robustness for different QA types. These results lead to an interesting preliminary conclusion: smaller models may overfit to the single-choice task format, while larger models seem to have learned more fundamental compositional spatial intelligence.
Table 3: Performance of MLLMs on single-choice and multiple-choice QA pairs. Results show that existing models, especially smaller ones, tend to overfit to single-choice questions.
[…]
InternVL2.5 Series
InternVL2.5-1B internvl2.5 32.4 22.2 38.0 22.6 28.3 2.0 7.4 7.4 5.9 5.4 28.7 5.6 0.20
InternVL2.5-2B internvl2.5 34.8 29.9 39.8 28.8 17.2 8.9 0.0 8.9 3.5 0.0 30.1 4.2 0.14
InternVL2.5-8B internvl2.5 14.1 48.1 55.4 33.6 34.5 13.9 23.2 33.5 13.9 11.3 37.3 19.1 0.51
InternVL2.5-38B internvl2.5 58.2 52.7 36.8 31.6 61.9 19.8 46.3 33.5 17.3 43.1 48.2 32.0 0.66
InternVL2.5-78B internvl2.5 54.6 58.8 61.7 40.3 55.8 17.3 35.5 21.3 37.4 50.6 54.2 32.4 0.60
Cambrian Series
Cambrian-8B cambrian 26.7 34.8 32.6 32.3 41.4 0.5 5.9 12.8 5.9 20.6 33.5 9.1 0.27
Cambrian-13B cambrian 16.0 27.7 32.8 30.5 31.0 0.5 2.0 10.3 6.4 15.0 27.6 6.8 0.25
Cambrian-34B cambrian 23.4 31.9 27.0 24.7 23.2 0.0 3.9 6.8 6.4 2.5 26.0 3.9 0.15
Others
VILA1.5-8B vila 15.5 37.0 43.2 24.1 26.7 0.0 0.5 1.5 2.0 26.9 29.3 6.1 0.21
Table 4: Accuracy comparison on basic and compositional QA pairs. The results reveal the relationship between the performance and integration of capacities.
Task Integrated Capability InternVL2.5-1B 8B Qwen2.5-VL-3B 7B Overall-C(%)
SA $\times$× C1,C2,C4,C8 37.8 66.4 60.0 64.8 57.3
$\checkmark$ +C7 22.2 ($\downarrow$ 15.6) 48.1 ($\downarrow$ 18.3) 46.7 ($\downarrow$ 13.3) 49.3 ($\downarrow$ 15.5) 41.6 ($\downarrow$ 15.7)
$O$ $\times$× C1,C2,C3,C8,C9,C10 52.0 66.4 56.4 41.2 54.0
$\checkmark$ +C7 38.0 ($\downarrow$ 14.0) 55.4 ($\downarrow$ 11.0) 50.7 ($\downarrow$ 5.7) 37.3 ($\downarrow$ 3.9) 45.4 ($\downarrow$ 9.6)
OS $\times$× C1,C2,C3,C9,C10 44.2 54.8 50.8 54.3 51.0
$\checkmark$ +C7 22.6 ($\downarrow$ 21.6) 33.6 ($\downarrow$ 21.2) 31.6 ($\downarrow$ 19.2) 32.3 ($\downarrow$ 22.0) 30.0 ($\downarrow$ 21.0)
EP $\times$× C1,C2,C8 66.6 75.8 63.4 65.3 67.8
$\checkmark$ +C7,C9 32.4 ($\downarrow$ 34.2) 14.1 ($\downarrow$ 61.7) 11.0 ($\downarrow$ 52.4) 15.0 ($\downarrow$ 50.3) 18.6 ($\downarrow$ 49.2)
FR $\times$× C1,C2,C3,C6,C8 70.9 89.7 85.6 85.7 83.0
$\checkmark$ +C7,C9,C10 28.3 ($\downarrow$ 42.6) 34.5 ($\downarrow$ 55.2) 25.3 ($\downarrow$ 60.3) 24.7 ($\downarrow$ 61.0) 28.2 ($\downarrow$ 54.8)
4.3 Capability Analysis
Atomic Capability vs. Compositional Capabilities. Firstly, the Tab. 4 shows that on the 5 types of questions with more compositional abilities, the performance of the four models all drops significantly. For the 3 questions with integrating C7 capability, the models’ accuracy in SA, OO, and OS tasks decreases by 15.7%, 9.6% and 21.0%, respectively. Despite the drop, the accuracy of SA and OO still remains at a decent level. Secondly, in the EP task with both C7 and C9 abilities integrated, the models’ overall accuracy plummets from 67.8% without integration to 18.6%, a huge drop of 49.2%. Similarly, for the FR task, which with the most compositional capacities, the four models’ average score crashes from 83.0% to 28.2%, an even larger drop of 54.8%.
These results partially reveal the relationship between accuracy and capability. As more abilities are incorporated, model performance declines to varying degrees, with some decreases even exceeding 60%. This indicates that current models have a limited grasp of integrated spatial intelligence, thereby highlighting the necessity of SpaCE-10.
Spatial Capability Breakdown. To better understand the strengths and weaknesses of the current MLLMs across various atomic capabilities, we constructed a model accuracy and atomic capability score matrix to associate QA accuracy with capabilities. As shown in Fig. 4, from the results, we observe several findings. Overall, different models show similar trends in capabilities such as C1 (Object Recognition), C2 (Spatial Localization), C4 (Size Comparison), C6 (Functional Understanding), and C9 (Reverse Reasoning), with scores in these areas predominantly blue, indicating stable performance across models. Capabilities like C5 (Counting) show green scores, indicating weaker performance across models. Notably, the distribution of capabilities across models is relatively consistent, suggesting that while these capabilities are not fully orthogonal, models have a similar overall grasp of different abilities.
Figure 4: Results of 8 representative MLLMs on 10 atomic capabilities of SpaCE-10. Each value reflects the model’s average accuracy (%) across all question types involving the respective spatial capability (C1–C10), as defined in the benchmark’s task-to-capability mapping.
Based on the macro analysis in the first step, we can now focus on the scores of each model. Similar to the accuracy trend, LLaVA-OneVision-72B demonstrates the most powerful performance, achieving the highest scores in over seven atomic capabilities. Among all its abilities, the one with the smallest gap to human performance is C4, with only a 14.2% difference from humans. The next strongest ability is C9, where the model’s performance in reverse reasoning is only 18.8% behind humans. However, GPT-4o performs the worst in C9, scoring only 32.3%, a 38.5% gap from humans. For most models, C4 is the best mastered atomic capability, indicating that current models have relatively strong capabilities in size comparison. However, in C5, counting abilities, all models perform poorly, with the highest score reaching only 38.8%, a 33.2% gap from humans. Therefore, MLLMs’ counting ability needs significant improvement.
Additionally, in terms of average scores, the best model, LLaVA-OneVision-72B, only achieved 47.2%, with a 24.2% gap from humans. The best closed-source model, GPT-4o, even has a 34.1% gap from humans. These results indicate that there is still a large room for improvement in both individual atomic spatial capabilities and integrated spatial intelligence in current MLLMs. This capability analysis provides valuable insights for guiding future enhancements of MLLMs. More experiments are attached in the supplementary materials.
5 Broader Impacts and Limitations
Although SpaCE-10 has undergone rigorous manual review, there may still be some annotation errors that slightly affect the evaluation of the model. In addition, the visual content of SpaCE-10 all comes from public datasets, which may contain some unreliable content. Besides, SpaCE-10 mainly focuses on indoor scenes, and the scene scale is relatively small, which means that the conclusions of SpaCE-10 may not be applicable to outdoor or large-scale scenes.
6 Conclusion
In this paper, we propose SpaCE-10, a comprehensive benchmark for evaluating compositional spatial intelligence in Multimodal Large Language Models (MLLMs). SpaCE-10 covers evaluations of MLLMs from 10 atomic spatial capabilities to 8 compositional capabilities. In SpaCE-10, we collect images and point clouds from 800 scenes and design a hierarchical annotation pipeline to produce over 5k high-quality question-answer pairs, covering various evaluation settings of MLLMs. To the best of our knowledge, SpaCE-10 is the most comprehensive benchmark for compositional spatial intelligence. Through extensive evaluation, we reveal critical limitations of current MLLMs and draw several findings that are beneficial to future work in the community. We believe these studies will provide an invaluable hint for future research toward human-level machine intelligence.
Appendix A More Detailed Illustration
A.1 QA Definition
In this section, we introduce the definition of each QA type and show the QA examples in Fig. 5. Notably, each QA type is the integration of multiple capabilities, and we also show the mapping between QA and capabilities in this figure and Tab. 5. The following are definitions:
Entity Quantification (EQ): Counting the number of objects in a scene.
Scene Quantification (SQ): Counting the number of regions within a scene.
Size Assessment (SA): Comparing the size relationships between different objects.
Object-Object Spatial Relationship (OO): Understanding the relative spatial relationship between two objects.
Object-Scene Spatial Relationship (OS): Understanding the relative spatial relationship between an object and the overall scene.
Functional Reasoning (FR): Reasoning objects that match or do not match certain functions based on relative spatial relationships.
Spatial Planning (SP): Global spatial navigation and path planning.
Figure 5: Examples of all types of QA. The blue examples represent the perception QA, and the purple ones mean the reasoning QA. The green circles are the correct answer.
A.2 Capability Definition
In this section, we illustrate the definition of each capability:
C1-Object Recognition: Identify what the object is.
C2-Spatial Location: Localizing an object in space.
C3-Spatial Relationship: Understanding relative spatial position relationship.
C4-Size Comparison: Comparing the size relationship of objects.
C5-Counting: Count the number of objects and scenes.
C6-Function Knowledge: Understanding the function of objects.
C7-Multi-view Fusion: Understanding spatial information from multiple views
C8-Forward Thinking: Understand the forward instructions of user instructions and execute them.
C9-Reverse Reasoning: Understand the reverse instructions of user instructions and execute them.
C10-Situated Observation: Imagine standing in a designated position in space and observing and understanding the scene.
Appendix B More Experiments
B.1 More Mulitple-Choice Results
In this section, we show the single-choice and multiple-choice results of more MLLMs in Tab. 6. As we mentioned in the main paper, the multiple/single score reveals a clear trend: most small-scale models exhibit notably low scores, indicating a strong tendency to overfit to single-choice questions. This behavior reflects a reliance on shallow pattern recognition rather than a true capability for handling more compositional spatial understanding tasks. In contrast, larger models demonstrate significantly improved performance on this metric, suggesting their enhanced ability to reason over complex spatial questions.
Table 5: Mapping of QA types and Spatial Capabilities. Each type of QA is the integration of multiple capabilities.
Tasks c1 c2 c3 c4 c5 c6 c7 c8 c9 c10
Entity Quantification (EQ) $\checkmark$ $\checkmark$ - - $\checkmark$ - $\checkmark$ $\checkmark$ - -
Scene Quantification (SQ) $\checkmark$ - - - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ - -
Size Assessment (SA) $\checkmark$ $\checkmark$ - $\checkmark$ - - $\checkmark$ $\checkmark$ - -
OO-Spatial Relationship (OO) $\checkmark$ $\checkmark$ $\checkmark$ - - - $\checkmark$ $\checkmark$ - $\checkmark$
OS-Spatial Relationship (OS) $\checkmark$ $\checkmark$ $\checkmark$ - - - $\checkmark$ $\checkmark$ - $\checkmark$
Entity Presence (EP) $\checkmark$ $\checkmark$ - - - - $\checkmark$ $\checkmark$ $\checkmark$ -
Functional Reasoning (FR) $\checkmark$ $\checkmark$ $\checkmark$ - - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Spatial Planning (SP) $\checkmark$ $\checkmark$ $\checkmark$ - - - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 6: More experiment results of multiple choices.
[…]
3D MLLMs
GPT4Scene-Qwen2-VL-7B gpt4scene 24.5 42.9 41.6 32.3 28.0 0.0 0.0 1.5 2.5 26.9 33.8 6.1 0.18
Close Source MLLMs
GPT-4o-2024-11-20 gpt4 51.7 63.9 63.2 41.6 37.0 20.3 43.3 40.9 18.8 31.9 51.4 31.0 0.60
Gemini-2.0-flash-thinking gemini 42.2 57.1 40.8 25.5 36.7 41.1 23.2 29.6 6.9 30.6 44.3 26.2 0.59
Claude-3.7-sonnet-20250219 claude3 26.0 44.4 39.3 26.8 45.0 24.3 30.5 31.4 9.4 35.0 36.3 26.1 0.72
Open Source MLLMs
InternVL2.5-1B internvl2.5 32.4 22.2 38.0 22.6 28.3 2.0 7.4 7.4 5.9 5.4 28.7 5.6 0.20
InternVL2.5-2B internvl2.5 34.8 29.9 39.8 28.8 17.2 8.9 0.0 8.9 3.5 0.0 30.1 4.2 0.14
Qwen2.5-VL-3B-Instruct qwen2.5 11.0 46.7 50.7 31.6 25.3 13.9 17.2 20.2 11.4 13.1 33.0 15.1 0.46
VILA1.5-3B vila 12.9 31.7 34.6 31.6 35.3 0.0 1.0 2.5 2.0 19.4 29.2 5.1 0.18
InternVL2.5-4B internvl2.5 36.0 50.2 50.7 16.2 51.7 11.9 25.1 25.6 17.3 43.8 40.9 24.7 0.60
Qwen2.5-VL-7B-Instruct qwen2.5 15.0 49.3 37.3 32.3 24.8 17.3 18.7 16.7 17.3 18.8 31.7 17.7 0.56
LLaVA-v1.5-7B liu2023visual 13.3 39.6 44.2 30.1 63.0 0.6 4.9 10.3 6.9 35.0 38.0 11.5 0.30
LLaVA-OneVision-7B llava-onevision 13.1 50.0 50.6 34.5 52.2 4.0 22.7 26.1 10.9 19.4 40.0 16.6 0.41
InternLM-XComposer2.5-7B internlm-xcomposer 6.6 39.4 41.5 34.9 29.8 1.98 22.1 25.6 16.3 20.0 30.4 17.1 0.56
Cambrian-8B cambrian 26.7 34.8 32.6 32.3 41.4 0.5 5.9 12.8 5.9 20.6 33.5 9.1 0.27
VILA1.5-8B vila 15.5 37.0 43.2 24.1 26.7 0.0 0.5 1.5 2.0 26.9 29.3 6.1 0.21
InternVL2.5-8B internvl2.5 14.1 48.1 55.4 33.6 34.5 13.9 23.2 33.5 13.9 11.3 37.3 19.1 0.51
Cambrian-13B cambrian 16.0 27.7 32.8 30.5 31.0 0.5 2.0 10.3 6.4 15.0 27.6 6.8 0.25
CogVLM2-LLaMA3-Chat-19B cogvlm 47.2 43.2 49.3 33.6 64.5 11.8 21.6 22.6 10.3 24.3 47.5 18.1 0.38
Cambrian-34B cambrian 23.4 31.9 27.0 24.7 23.2 0.0 3.9 6.8 6.4 2.5 26.0 3.9 0.15
InternVL2.5-26B internvl2.5 56.7 54.9 54.0 33.0 37.3 29.7 28.1 32.5 17.8 31.3 47.1 27.8 0.59
Qwen2.5-VL-32B-Instruct qwen2.5 11.7 51.3 38.1 32.3 11.4 11.7 23.6 14.8 7.4 42.5 29.1 20.0 0.69
InternVL2.5-38B internvl2.5 58.2 52.7 36.8 31.6 61.9 19.8 46.3 33.5 17.3 43.1 48.2 32.0 0.66
VILA1.5-40B vila 10.1 52.0 58.4 40.9 31.4 11.9 35.5 37.4 19.3 24.3 38.5 25.6 0.67
LLaVA-OneVision-72B llava-onevision 66.6 64.1 62.5 40.3 65.9 34.2 34.5 32.5 16.3 35.0 59.8 30.5 0.51
Qwen2.5-VL-72B-Instruct qwen2.5 26.0 53.7 37.8 32.1 31.1 14.5 27.6 17.7 13.4 30.0 36.1 20.6 0.57
InternVL2.5-78B internvl2.5 54.6 58.8 61.7 40.3 55.8 17.3 35.5 21.3 37.4 50.6 54.2 32.4 0.60
Among the close-source MLLMs, Claude-3.7-sonnet achieves the highest score with 0.72, showing its strong ability in both single and multiple task. For the open-source MLLMs, the Qwen series stands out, with Qwen2.5-VL-32B achieving the second highest score of 0.69, showcasing strong competence in processing multi-option (double-choice) tasks that require integrative spatial reasoning. The InternVL series also performs consistently well: all larger variants (4B, 8B, 26B, 38B, 78B) surpass the 0.5 threshold, excluding only the 1B and 2B versions. These results highlight the structural advantages and training effectiveness of these models when facing combinatorial challenges. Furthermore, all close-source models (GPT-4o, Gemini, Claude) score around 0.6, demonstrating their overall strength in tasks demanding structured spatial comprehension.
On the other hand, several models yield unexpectedly poor results. For instance, Cambrian-34B scores only 0.15—almost identical to the much smaller InternVL2.5-2B (0.14)—suggesting a significant deficiency in understanding spatial compositions and relationships. Similarly, although LLaVA-OneVision-72B achieves state-of-the-art performance in overall metrics, its multiple/single score of just 0.51 falls behind comparably sized models such as Qwen2.5-VL-72B (0.57) and InternVL2.5-78B (0.60), indicating a lingering tendency to overfit to single-choice formats.
These findings underscore the importance of addressing task format overfitting in future MLLMs training. They also reveal current limitations in compositional spatial intelligence, validating the value and potential of the proposed SpaCE-10 benchmark. By highlighting nuanced weaknesses in compositional spatial comprehension across diverse models, this benchmark serves as a crucial diagnostic tool for future spatial intelligence evaluation of MLLMs.
B.2 Pipeline Component Ablation
To systematically evaluate the contribution of each key component in the structured data generation pipeline to the accuracy of basic QA generation, we designed a stepwise ablation study. Specifically, we selected five QA task types (SA, OO, OS, EP, FP), and for each type, we randomly sampled 30 scenes, generating one question per scene. The accuracy of each generated question was manually verified by human experts.
The experimental results clearly demonstrate the cumulative gains of each module. Using the 2D Captioner alone as the baseline, the generated QA already showed relatively stable accuracy across most categories (an average of 62.0%), with particularly high accuracy in the EP task at 96.7%. This reflects the relatively low difficulty of generating questions for this category and that 2D visual information sufficiently supports it. With the addition of the 3D Captioner, the overall accuracy improved significantly (an average increase of 14.7%), indicating that 3D information effectively supplements the limitations of 2D vision and enhances the model’s understanding of spatial and object attributes. Further incorporating the Inspector component led to another increase in accuracy, reaching 79.3% (a 17.3% improvement over the baseline), showing that this module plays an important role in validating and refining the details of question generation. Finally, after adding structured data, the overall average accuracy reached 84.0%, a 22% improvement compared to the initial baseline, fully demonstrating the critical value of structured information in improving the quality of basic QA generation.
Table 7: Ablation study of structural data generation pipeline. We randomly sample 30 scenes for each type of QA and generate one QA for each scene. The results demonstrate the effectiveness of each component.
Component SA $O$ OS EP FP Average
2D Captioner 50.0 50.0 46.7 96.7 66.7 62.0
+ 3D Captioner 80.0 63.3 63.3 96.7 80.0 76.7 (+14.7)
+ Inspector 80.0 66.7 70.0 96.7 83.3 79.3 (+17.3)
+ Structural Data 86.7 76.7 76.7 96.7 83.3 84.0 (+22.0)
Appendix C Case Study
C.1 Basic QA Quality
In this paper, we have developed a sophisticated pipeline for the automated generation of high-quality Basic QA (Basic Question Answering). In the previous section, we systematically verified the effectiveness of each component within the pipeline. This section presents two typical high-quality QA cases and delves into them with specific context-based analysis and discussion.
As shown in Fig. 6, the first case centers on the size estimation task (Basic SA). As depicted in the figure, through meticulously crafted prompts and high-quality snapshots, we steered GPT to accurately generate questions. When comparing the volume relationship between a white wooden baby crib and a wardrobe, GPT delivered an impressive performance. It not only correctly identified the height and width dimensions of each object but also precisely determined their relative differences across multiple perspectives. For instance, its response accurately pointed out that the wardrobe is taller than the crib but has a similar width, demonstrating a good grasp of the 3D geometric properties of objects.
The second case focuses on the spatial relationship understanding task (Basic OO), as shown in the Fig. 7, in the context of judging the spatial relationship between a yellow chair and an entrance. In this example, GPT not only accurately distinguished the relative direction of "left" but also correctly identified the spatial distance difference between "near" and "far". Although the question only involved one yellow chair, the distractors in the options were somewhat deceptive, which in turn enhanced the question’s ability to assess the model’s spatial understanding. This case also indirectly confirms the feasibility and rationality of using GPT for generating spatial intelligence QA questions. Through these two cases, we observed that powerful MLLMs, when only provided with high-quality 2D visual inputs, are already capable of understanding certain 3D spatial information, such as object size, volume relationships, and spatial orientation. This ability suggests a promising path for future exploration: unlike current 3D models that sacrifice conversational abilities to fit point cloud data, 2D MLLMs can still demonstrate strong spatial understanding potential even without explicitly incorporating 3D structural modeling.
Figure 6: Showcase of Basic SA QA. ‘H’ and ‘W’ mean the height and width of objects, respectively. In this case, GPT-4o shows the precise understanding of both size and geometry.
Figure 7: Showcase of Basic $O$ QA. This case reflects the capability in spatial relationships and QA design. Notably, all spatial relationship is based on the observation at the entrance.
Figure 8: Interface of annotation tools.
C.2 Annotation Interface
To facilitate efficient annotation, we developed a custom annotation tool, with the interface shown in the Fig. 8. During the annotation process, human experts are restricted to viewing only 3D snapshots to judge the correctness of each QA pair—they are not allowed to view the 2D images. This design offers two main advantages: (1) It ensures a high annotation speed. By examining only a small number of 3D snapshots, annotators can quickly grasp the overall layout of the scene while significantly reducing their visual workload. (2) Since 2D images contain overly fine-grained details—many of which may not be present in the 3D scene—using only 3D information to filter out incorrect questions helps ensure that the resulting QA pairs are suitable for evaluation across both 2D and 3D models. Additionally, during evaluation, we tag erroneous questions to ensure none are overlooked. This end-to-end process not only prioritizes annotation efficiency but also reflects our rigorous commitment to data quality control.